Read the paper at your pace. Open an explanation when you need it.
Mathematics (120 questions)
Question 1
Suppose ω is a cube root of unity with ω ≠ 1. Suppose P and Q are the points on the complex plane defined by ω and ω2. If O is the origin, then what is the angle between OP and OQ?
2. Symmetric: If any one element is related to any other element, then the second element is related to the first.
R is symmetric if for all x, y ∈ A, if xRy, then yRx.
<br>
3. Transitive: If any one element is related to a second and that second element is related to a third, then the first element is related to the third.
R is transitive if for all x, y, z ∈ A, if xRy and yRz, then xRz.
<br>
Calculation:
Given, x, y and z are positive numbers i.e 1, 2, 3 and so on
1. Reflexive:
To show x ≤ x2
x2 – x ≥ 0
x ( x – 1 ) ≥ 0
So, for x ≥ 1 or x ≤ 0, the above relation is true. It is given that the x > 0 so the acceptable value will be,
x ≥ 1
So, for all x ≥ 1, the given relation will be reflexive but for the value of xϵ (0, 1), it will not be reflexive.
So, the relation is not reflexive for all the positive values of x.
2. Symmetric:
To show when x ≤ y2 leads to y ≤ x2
Considering x = 1 and y = 2
x2 = 1 and y2 = 4
Here x ≤ y2 as 1 ≤ 4 but y ≤ x2 is not true as y = 2 is greater than x2 = 1
Thus relation * is not symmetric
3. Transitive: To show when xRy or x ≤ y2 and yRz or y ≤ z2 leads to xRz or x ≤ z2
Considering x = 5, y = 3 and z = 2
⇒ x2 = 25 , y2 = 9 and z2 = 4
Here x ≤ y2 as 5 ≤ 9, y ≤ z2 as 3 ≤ 4, but x ≤ z2 is not true as x = 5 is greater than z2 = 4
Thus relation * is not transitive
Question 3
If x2 - px + 4 > 0 for all real values of x, then which one of the following is correct?
The term b2- 4ac is called discriminant, it is important as it tells about the nature of roots.
1. ;b2−4ac<0
There are no real roots
1. ;b2−4ac=0
There is one real root
1. ;b2−4ac>0
There are two real roots
A parabola graphically represents a quadratic equation with a vertex located at the origin, below the x-axis, or above the x-axis. Therefore, a quadratic function may have one, two, or zero roots.
Calculation:
Given: x2;−;px;+;4;>;0;;∀;real;values;of;x
On comparing the given equation with the standard quadratic equation
ax2+bx;+;c=;0, we get
a = 1, b = -p and c = 4
Given quadratic equation is more than zero that means it has no real roots If roots are not real (Then only the quadratic equation will not intersect the x-axis)
∴ b2−4ac<0
⇒(−p)2−4(1)(4)<0
⇒(p)2−16<0
⇒(p)2−42<0
⇒(p+4)(p−4)<0 (Using a2- b2 = (a + b). (a - b))
This inequality holds if p ranges from -4 to 4 i.e., p = (-4, 4)
Only |p| < 4 satisfy the above condition
Hence, option 1 is correct.
Important Points
If the quadratic function is always more than zero, that means, when graphed on the coordinate system, the function never intersects the X-axis. That means that the quadratic equation has no real roots.
Question 4
If z=x+iy=(21−2i)−25, where i=−1, then what is the fundamental amplitude of z−i2z−2?
Let A(x1, y1) and B(x2, y2) be any two points on the X-Y plane. Suppose point C is the mid-point of the line segment AB, then the co-ordinates of point C is:
Midpoint;=;(2x1;+;;x2,;2y1;+;;y2)
<br>
Calculation:
Let the intersection on x axis (P) be (x, 0)
Let the intersection on y axis (Q) be (0, y)
Given mid-point of PQ: (3 , 5 )
Using midpoint formula:
20;+;x=3
∴ x = 6
20;+;4=5
∴ y = 10
So, P (6, 0) and Q (0, 10)
Since the triangle is formed with sides as both the coordinate axes it is right angled at O.
PO = 6
QO = 10
Area=21×PO×QO=26:×:10;=30
Question 8
If a circle of radius b units with center at (0, b) touches the line y = x -√2, then what is the value of b?
Greatest Integer Function: Greatest Integer Function [x] indicates an integral part of the real number x which is nearest and smaller integer to x. It is also known as floor of x
In general, If n≤ x ≤ n+1 Then [x] = n (n ∈ Integer)
Means if x lies in [n, n+1) then the Greatest Integer Function of x will be n.
If a sequence is given in the form of Arithematic progression (A.P.) and all its terms are equal, i.e.
a, a, a, . . .
then such A.P. are called as Constant A.P. and has a common difference = 0.
\( \smallint sin2xdx=-\frac{cos~2x}{2}\)
Calculation:
We are given the sequence \(a_n= \mathop \smallint \limits_0^{\rm{\pi }} \frac{{{{\sin }^2}\left{ {\left( {{\rm{n}} + 1} \right){\rm{x}}} \right}}}{{\sin 2{\rm{x}}}}{\rm{dx}}\)
Area of a triangle whose vertices are (x1, y1), (x2, y2) and (x3, y3), is given by the expression Area = (1/2) [x1 (y2- y3) + x2 (y3-y1) + x3(y1-y2)]
Whose expression can be written in the form of a determinant as: \({\rm{Area}} = {\rm{;}}\frac{1}{2}\left| {\begin{array}{*{20}{c}} {{{\rm{x}}_1}}&{{{\rm{y}}_1}}&1\ {{{\rm{x}}_2}}&{{{\rm{y}}_2}}&1\ {{{\rm{x}}_3}}&{{{\rm{y}}_3}}&1 \end{array}} \right|\)
Calculation:
Given lines are y = 3x, y = 6x and y = 9
Let,
Side AB represent line y = 3x,
Side AC represent line y = 9 and
Side BC represent line y = 6x
So for coordinates of point A, we will consider sides AB, AC and then solve respective equation of lines
We have,
y = 3x and y = 9
⇒ 3x = 9
∴ x = 3
Therefore A = (3, 9)
Similarly, for point B,
We have,
y = 3x and y = 6x
⇒ 6x = 3x
∴ x = 0
Therefore B = (0, 0)
An also, for point C,
We have,
y = 6x and y = 9
⇒ 6x = 9
∴ x = 3/2
Therefore C = (3/2, 9)
As we know that, \({\rm{Area}} = {\rm{;}}\frac{1}{2}\left| {\begin{array}{*{20}{c}} {{{\rm{x}}_1}}&{{{\rm{y}}_1}}&1\ {{{\rm{x}}_2}}&{{{\rm{y}}_2}}&1\ {{{\rm{x}}_3}}&{{{\rm{y}}_3}}&1 \end{array}} \right|\)
For local maxima: Sign of f’(x) changed from positive to negative.
So, Local minima at x = 1
Hence, there are one local maxima.
Question 31
Let f(x) and g(x) be twice differentiable functions on [0, 2] satisfying f’’(x) = g”(x), f’(1) = 4, g’(1) = 6, f(2) = 3 and g(2) = 9. Then what is f(x) – g(x) at x = 4 equal to?
\({\rm{f}}\left( {\rm{x}} \right) = \left| {\begin{array}{*{20}{c}} {{{\rm{x}}^3}}&{\sin {\rm{x}}}&{\cos {\rm{x}}}\ 6&{ - 1}&0\ {\rm{p}}&{{{\rm{p}}^2}}&{{{\rm{p}}^3}} \end{array}} \right|\), where p is a constant
Differentiation of the determinant of a matrix: Differentiate the first row (or column) and keep the entries of the other row untouched. Do this for all other rows (or columns). Then find the sum of all such determinants. Then you can get a determinant.
\({\rm{f}}\left( {\rm{x}} \right) = \left| {\begin{array}{*{20}{c}} {{{\rm{x}}^3}}&{\sin {\rm{x}}}&{\cos {\rm{x}}}\ 6&{ - 1}&0\ {\rm{p}}&{{{\rm{p}}^2}}&{{{\rm{p}}^3}} \end{array}} \right|\), where p is a constant
Greatest Integer Function: Greatest Integer Function [x] indicates an integral part of the real number x which is nearest and smaller integer to x. It is also known as floor of x
In general, If n ≤ x ≤ n + 1 Then [x] = n (n ∈ Integer)
Means if x lies in [n, n + 1) then the Greatest Integer Function of x will be n.
A parallelogram is a quadrilateral with two pairs of parallel sides. The opposite sides of a parallelogram are equal in length and the opposite angles are equal in measure.
Diagonals of a parallelogram bisect each other.
Let A(x1, y1) and B(x2, y2) be any two points on the X-Y plane. Suppose point C(x, y) is the mid-point of the line segment AB, then the coordinates of point C is: (x,y)=(2x1;+x2,;2y1+;y2)
Let A(x1, y1) and B(x2, y2) be any two points on the X-Y plane. Distance;between;A;and;B;=;(x2−;x1)2+;(y2−;y1)2
Calculation:
Here, A (1, 2), B (4, y), C (x, 6) and D (3, 5) are the vertices of a parallelogram.
AB = CD and AD = BC (sides of parallelogram)
AC and BD are the diagonals which bisect each other.
Suppose diagonals intersect each other at (m, n)
Using midpoint formula, we get
Midpoint of AC = (21+x,26+2)
Midpoint of BD = (24+3,2y+5)
We know that diagonals of a parallelogram bisect each other.
∴ Midpoint of AC = midpoint of BD
(21+x,26+2)=(24+3,2y+5)
∴21+x=27;and;28=2y+5
⇒1+x=7;and;8=y+5
⇒x=6;and;y=3
Now, Coordinates of B (4, y) = B (4, 3) and Coordinates C (x, 6) = (6, 6)
Let's find the distance,
AC=;(6−1)2+(6−2)2=;52+42=;25+16=41
⇒ AC2 = 41
BD=;(3−4)2+(5−3)2=;−12+22=;1+4=5
⇒ BD2 = 5
Now,
AC2 – BD2 = 41 - 5 = 36
Hence option 3 is the correct answer.
Question 59
What is the point of intersection of the diagonals?
Greatest Integer Function: Greatest Integer Function [x] indicates an integral part of the real number x which is nearest and smaller integer to x. It is also known as floor of x
In general, If n≤x;≤n+1; Then [x] = n (n ∈ Integer)
Means if x lies in [n, n+1) then the Greatest Integer Function of x will be n.
<br>
Function modulus:
|f(x)| = f(x) whenever f(x) ≥ 0
|f(x)| = -f(x) whenever f(x) < 0
Calculation:
Given: f(x) be the greatest integer function and g(x) be the modulus function.
Order: The order of a differential equation is the order of the highest derivative appearing in it.
Degree: The degree of a differential equation is the power of the highest derivative occurring in it, after the Equation has been expressed in a form free from radicals as far as the derivatives are concerned.
<br>
Calculation:
Given:
y = cx + c2 – 3c3/2 + 2 …. (1)
Differentiate with respect to x, we get
⇒ dy/dx = c
Substitute the value of c in equation 1st, we get
⇒y=xdxdy+;(dxdy)2−3(dxdy)23+2
⇒3(dxdy)23=(dxdy)2+xdxdy−y+2
Squaring both sides, we get
⇒9(dxdy)3=(dxdy)4+…
Hence order is 1 and degree is 4.
Question 79
What is \(\mathop \smallint \limits_{ - 2}^2 {\rm{x;dx}} - \mathop \smallint \limits_{ - 2}^2 \left[ {\rm{x}} \right]{\rm{dx}}\) equal to, where [⋅] is the greatest integer function?
If f(x) is periodic function with period T then \(\mathop \smallint \nolimits_0^{{\rm{nT}}} {\rm{f}}\left( {\rm{x}} \right){\rm{dx}} = {\rm{n}}\mathop \smallint \nolimits_0^{\rm{T}} {\rm{f}}\left( {\rm{x}} \right){\rm{dx}},{\rm{;n}} \in {\rm{Z}}\)
Let A(x1, y1) and B(x2, y2) be any two points on the X-Y plane. Suppose point C is the mid-point of the line segment AB, then the coordinates of point C is:
(2x1+x2,2y1+y2).
Calculation:
Mid-point of (10, -6) and (k, 4) :
(a,2b)=(210+k,2−6+4)=(5+2k,−1)
Therefore,
a=2k+5
2b = - 1
Given, a – 2b = 7
(2k+5;)+1=7
2k+6=7
2k=1
k = 2
Question 89
Consider the following statements:
If ABC is an equilateral triangle, then 3 tan (A + B) tan C = 1
If ABC is a triangle in which A = 78°, B = 66°, then tan(2A+C)<tanA
If ABC is any triangle, then tan(2A+B)sin(2C)<cos(2C)
Mean deviation about the mean = \(\frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} \left| {{{\rm{x}}_{\rm{i}}}-{\rm{;\bar x}}} \right|}}{{\rm{n}}}\), Where xˉ is mean.
Calculation:
Given numbers are 10, 9, 21, 16, 24
Total numbers = 5
Mean=xˉ=;510+9+21+16+24=580=16
We know that mean deviation about the mean = \(\frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} \left| {{{\rm{x}}_{\rm{i}}}-{\rm{;\bar x}}} \right|}}{{\rm{n}}}\)
Mean deviation from the mean = 5∣10−;16∣+∣9−;16∣+∣21−;16∣+∣16;−;16∣+∣24−;16∣;
=56+7+5+0+8=526=5.2
Question 92
Three dice are thrown simultaneously. What is the probability that the sum on the three faces is at least 5?
Complement of an event: The complement of an event is the subset of outcomes in the sample space that are not in the event. A's complement is itself an event.
The probability of the complement of an event is one minus the probability of the event.
P (A’) = 1 - P (A), Where P(A) be the probability of Event A and P (A’) be probability of the complement of Event A
<br>
Calculation:
Given: A and B are two independent events.
P(A)=31 and P(B)=43
Probability that exactly one of the two events A or B occurs = P(A) P(B') + P(B) P(A')
Variance = \(\frac{{\sum {{\rm{x}}{\rm{i}}}^2}}{{\rm{N}}} - \frac{{{{\left( {\sum {{\rm{x}}{\rm{i}}}} \right)}^2}}}{{{{\rm{N}}^2}}}\), Where N is number of observation.
Probability of required event = 1 – (Probability of not required event)
<br>
Calculation:
To find:
The probability that the sum on the two faces is greater than or equal to 4
Let A be the event of getting the sum on the two faces is less than 4
∴ A’ = {(1,1), (1,2), (2,1)}
∴ n(A’) = 3
P(A′)=n(S)n(A′)=363=121
But we need the probability of the sum on the two faces is greater than or equal to 4
∵ Probability of required event = 1 – Probability of not required event)
Let, that event be A,
∴P(A) = 1 - P(A')
⇒P(A)=1−121
=1211
Hence, option (3) is correct.
Question 98
A certain type of missile hits the target with probability p = 0.3. What is the least number of missiles should be fired so that there is at least on 80% probability that the target is hit?
A point is chosen at random inside a rectangle measuring 6 inches by 5 inches. What is the probability that the randomly selected point is at least one inch from the edge of the rectangle?
Statement (1) is correct because total area of the rectangles in a histogram is equal to the total area bounded by the corresponding frequency polygon and x-axis.
Statement (2) is also correct because class intervals are unequal in a frequency distribution change into continuous class intervals.
Question 106
A fair coin is tossed 100 times. What is the probability of getting tails an odd number of times?
What is the number of ways in which 3 holiday travel tickets are to be given to 10 employees of an organization, if each employee is eligible for any one or more of the tickets?
If a=a1i^+a2j^+a3k^;and;b=b1i^+b2j^+b3k^ then \(\vec a \times ;\vec b = \left| {\begin{array}{*{20}{c}} {\hat i}&{\hat j}&{\hat k}\ {{a_1}}&{{a_2}}&{{a_3}}\ {{b_1}}&{{b_2}}&{{b_3}} \end{array}} \right|\)
If aand b are two vectors then the vector which is orthogonal to both the vectors a and b is given by:a×;b
If a is a vector then a unit vector in the direction of a is given by: a^=∣a∣a
Calculation:
Let a=i^+j^+k^ and b=2i^+3j^−k^.
As we know that, if a and b are two vectors then the vector which is orthogonal to both the vectors a and b is given by:a×;b
We also know that \(\vec a \times ;\vec b = \left| {\begin{array}{*{20}{c}} {\hat i}&{\hat j}&{\hat k}\ {{a_1}}&{{a_2}}&{{a_3}}\ {{b_1}}&{{b_2}}&{{b_3}} \end{array}} \right|\)
\(⇒ \vec a \times ;\vec b = \left| {\begin{array}{*{20}{c}} {\hat i}&{\hat j}&{\hat k}\ {{1}}&{{1}}&{{1}}\ {{2}}&{{3}}&{{-1}} \end{array}} \right| = - 4 \hat i + 3 \hat j + \hat k\)
Let c=−4i^+3j^+k^
As we know that if ais a vector then a unit vector in the direction of a is given by: a^=∣a∣a.
So, the unit vector in the direction of c is given by: c^=∣c∣c.
⇒∣c∣=26
Hence, unit vector along cis 26−4i^+3j^+k^
Question 111
If a,;b and c are the position vectors of the vertices of an equilateral triangle whose orthocentre is at the origin, then which one of the following is correct?
Elementary transformations does not changes the value of determinant of a matrix.
If any two rows (columns) of a matrix are same then the value of the determinant is zero
If all the elements of one row (column) is multiplied by the same quantity say k, then the value of new determinant is k times the value of original determinant.
If \(A = \left[ {\begin{array}{*{20}{c}} {{a_{11}}}&{{a_{12}}}&{{a_{13}}}\ {{a_{21}}}&{{a_{22}}}&{{a_{23}}}\ {{a_{31}}}&{{a_{32}}}&{{a_{33}}} \end{array}} \right]\) is a square matrix of order 3, then determinant of A is given by:
As we know that if all the elements of one row (column) is multiplied by the same quantity say k, then the value of new determinant is k times the value of original determinant.
\(\Rightarrow \left| {\begin{array}{{20}{c}} 1&a&{b + c}\ 1&b&{c + a}\ 1&c&{a + b} \end{array}} \right| = (a + b + c) \left| {\begin{array}{{20}{c}} 1&a&{1}\ 1&b&{1}\ 1&c&{1} \end{array}} \right|\)
As we know that, if any two rows (columns) of a matrix are same then the value of the determinant is zero.
Suppose a set of n objects has n1 of one kind of object, n2 of a second kind, n3 of a third kind, and so on with n = n1 + n2 + n3 +...+ nk then the number of distinguishable permutations of the n objects is = \(\frac{{{\rm{n}}!}}{{{{\rm{n}}_1}! \times {{\rm{n}}_2}! \times {{\rm{n}}3}! \ldots \ldots \ldots {{\rm{n}}{\rm{k}}}!}}\)
Calculation:
Given: Three 0’s and two 1’s
0
0
0
1
1
Total number = 3 + 2 = 5
In the set of 5 number has 3 number of one kind and 2 number of the second kind.
Therefore, Number of different messages that can be represented by three 0's and two 1's = 3!2!5!=10